# UniCTokens: Boosting Personalized Understanding and Generation via Unified Concept Tokens

### v2

@@ 4.3 Existing Personalized Understanding and Generation Benchmarks
In order to facilitate a fair comparison, we also evaluated our method on benchmarks for pure personalized understanding and generation. The results of the evaluation on Yo’LLaVA [36] and MC-LLaVA Datasets [24] are presented in Tab. 2 (left). Due to the limitations of the unified model’s capabilities on multi-concept tasks, we only conducted tests on the single-concept portion of MC-LLaVA. Our method outperforms the current leading unified personalized model, utilizing only 1.3 B parameters and fewer training images. Notably, our proposed UniCTokens outperforms on all understanding tasks with an average of 5.13% when compared to an important baseline, Yo’LLaVA(1.3B), demonstrating the potential of scaling UniCToken to achieve state-of-the-art performance.
We thoroughly evaluate personalized generation capabilities of UniCTokens on Dreambench [19] and Yo’LLaVA Dataset [36]. Our approach significantly enhances the capabilities of vanilla Show-o. Additionally, our performance surpasses that of Yo’Chameleon, particularly on CLIP-T, which measures the model’s ability for controllable generation. This improvement can be attributed to our more effective prompt design and mutual enhancement between personalized tasks. Notably, we also outperform the significant generative baseline, Dreambooth, with the same training data, demonstrating that our method can generate high-quality and realistic images containing user-provided concepts.
@@ 4.4 Ablation Study and Analysis
Different Training Strategies.
Our approach utilizes a three-stage training strategy to facilitate information transfer across tasks. We then validate this design, beginning from generation, especially examining the inclusion of Stage 1. Omitting Stage 1 yields a variant of Text inversion distinguished by an increasing parameter count. Tab. 3 emphasizes the essential role of Stage 1, particularly in personalized knowledge-driven generation, omitting Stage 1 leads to significant degradation in generation quality. This further corroborates the existence of information transfer across tasks in our design. Stage 3 does not update the generative parameters, resulting in a subtle influence on the generation task. In the context of the understanding task, we present the results of the evaluations conducted across all stages. Although performance in the understanding task declines in Stage 2 to facilitate generation, the subsequent injection of fine-grained visual information from Stage 3 enhances understanding capabilities. Results demonstrate that our approach achieves optimal performance, facilitating mutual enhancement between the two tasks.
Stage Personalized Understanding Personalized Generation Rec. VQA QA Pure Gen. People Gen. T2I Weight BLEU Score BLEU Score CLIP - I CLIP-T DINO Face-Simi Score CLIP-I Stage 1 0.637 0.494 0.538 0.540 0.600 0.524 0.278 0.446 0.060 0.204 0.511 Stage 2 w/o 1 0.616 0.479 0.488 0.527 0.581 0.695 0.243 0.582 0.210 0.297 0.695 Stage 2 0.621 0.497 0.532 0.538 0.605 0.752 0.280 0.648 0.349 0.349 0.750 Stage 3 0.790 0.503 0.523 0.544 0.603 0.750 0.282 0.646 0.334 0.334 0.749
@@ 5 Conclusion
We present UniCTokens, an innovative framework designed to personalize a unified VLM by training unified concept tokens. Our proposed three-stage training strategy enables UniCTokens to efficiently enhance personalized understanding and generation, facilitating cross-task information transfer to achieve mutual enhancement. Experimental results demonstrate that, while maintaining a smaller model size and fewer training samples, UniCTokens achieves state-of-the-art performance in concept understanding, concept generation, and knowledge-driven generation tasks. The insights derived from our analysis are noteworthy, shedding light on the development of unified VLMs. Furthermore, the advancement of personalized AI holds significant promise for improving human lives by enabling tailored interactions, enhancing creativity, and offering solutions that align with individual needs and preferences, thereby fostering a more intuitive and responsive technological ecosystem.
@@ Appendix A Additional Implement Detail
To clarify our three-stage training, we list all details as shown in Tab. 4.
@@ Appendix B Additional Experiment Setup
For understanding, we evaluate the model on three tasks: personalized recognition, VQA, and QA. In the Recognition task, we compute the recall for both positive and negative samples and report their arithmetic mean. For VQA and QA, we evaluate the predicted answers using BLEU [48] scores against ground truth, and further apply an LLM-based evaluation that scores responses from 0 to 1 based on alignment with key points in the reference answers. For general personalized image generation, we adopt prompts from the DreamBooth [19] dataset, and evaluate image quality using CLIP-based metrics: CLIP-I (image-to-image similarity) and CLIP-T (image-to-text similarity). Since half of our concepts are human subjects, we additionally employ the off-the-shelf ArcFace model [49] to measure facial similarity between generated and reference images. For the evaluated GPT-4o and GPT-4o used for scoring, we utilize different versions of them to make fair comparisons.
@@
  • Yo’LLaVA [36]: One of the earliest work to explore VLM personalization. Following the original paper, we manually construct hard negative datasets and train Yo’LLaVA with different sizes of base model(Phi-1.5 [50], 1.3B and Vicuna [51], 13B) to make a fair comparison.
  • MC-LLaVA [24]: A model designed for enhancing multi-concept personalization tasks. Utilizing dual textual and visual prompts, it serves as a strong baseline for personalized understanding.
  • RAP-MLLM [21]: We utilize the RAP-LLaVA model and follow the RAP-MLLM approach to construct a personalized database for each concept. To obtain the capability of understanding personalization, RAP-MLLM conducted a post-training on a dataset of 260K in size.
  • Dreambooth [19]: DreamBooth enhances the personalization capability of generative models by allowing users to fine-tune models with a limited number of images, thereby producing highly specific and context-aware outputs. For fair comparison, we utilize different number of training data (10, 3,000) to better evaluate Dreambooth.
  • Text Inversion [26]: Text inversion is a technique that transforms textual prompts into corresponding visual representations, enabling the generation of images based on description.
@@ Appendix C Catastrophe Forgetting
When a model acquires new knowledge, there exists the risk of catastrophic forgetting of previously learned information. Following Yo’Chameleon [22], we evaluate the personalized model on several benchmarks assessing general capabilities. We conducted a comparative analysis against the original Show-o across well-established multimodal benchmarks: GenEval [52], MMMU [53], and POPE [54]. The experimental results shown in Fig. 8 indicate that, despite training through a three-stage process and having task-specific tokens, model performance does not diminish after each stage, effectively preserving the model’s general capabilities.
@@
GenEval MMMU POPE
Original 0.68 26.7 80.0
Stage 1 0.67 26.6 80.0
Stage 2 0.66 26.6 80.0
Stage 3 0.66 26.6 80.0
Figure 7: Catastrophic Forgetting Evaluation. UniCTokens maintains performance similar to vanilla Show-o.
@@
Figure 8: Limitations. Example images under different style controls.
@@ Appendix D Additional Related Work
A multitude of efforts have been dedicated to employing a single model for both understanding and generation. SEED [55] adapted image representation through discrete tokenization for language modeling, leveraging autoregressive conditioning for generation via an external decoder.  [56, 57] restructured conditioning information aggregation but still relied on extra modules for generation. Chemeleon [30] is an early hybrid unified model capable of generating and understanding intert-leaved text-image content. Janus [28] claims that understanding and generation require distinct information, employing different tokenizers for each task. Emu3 [58] converts images, text, and video into discrete tokens, enabling joint training of a Transformer on multimodal sequences.  [46, 59] utilize a hybrid of autoregressive and diffusion methods for text and image processing. The above-mentioned work all focuses on general tasks, neglecting exploration in personalization scenarios. In this paper, we employ Show-o [46] as the backbone to efficiently achieve unified personalization without forgetting the general capability of the model.
@@
The establishment of unified models seeks to optimize the strengths of understanding and generation, allowing information transfer between tasks and mutual improvement. Early efforts to link these tasks mainly utilized a serial processing paradigm.  [60, 61, 62] employed image captioning models to generate textual conditions for text-to-image models. While these methods highlight the potential of leveraging understanding for generation, they lack end-to-end optimization and frequently suffer from information loss.  [63, 64, 65, 66, 67] explored deeper integration of understanding and generation. MetaQuery [68] utilizes learnable queries on frozen VLMs to extract conditions for generation, but it primarily emphasizes how understanding aids generation while neglecting the inverse. In this paper, we propose a three-stage training strategy that achieves information transfer and mutual enhancement between tasks in personalization scenarios, providing insights for the broader development of unified models.
@@ Appendix E Additional Qualitative Results
Appendix F Additional Ablation Study
@@
Learnable Token Length Und. Gen.
@@ Appendix G Details of Dataset
The dataset’s data sources comprise animals and objects obtained from MC-LLaVA [24], Yo’LLaVA [36] and MyVLM [23], with images of individuals sourced from Yo’LLaVA and various online platforms. All images and generated training data have been subjected to rigorous human validation processes. We present a set of images along with extra textual information in Fig. 11.

### v3

@@ 4.3 Existing Personalized Understanding and Generation Benchmarks
In order to facilitate a fair comparison, we also evaluated our method on benchmarks for pure personalized understanding and generation. The results of the evaluation on Yo’LLaVA nguyen2024yo and MC-LLaVA Datasets an2024mc are presented in Tab. 2 (left). Due to the limitations of the unified model’s capabilities on multi-concept tasks, we only conducted tests on the single-concept portion of MC-LLaVA. Our method outperforms the current leading unified personalized model, utilizing only 1.3 B parameters and fewer training images. Notably, UniCTokens outperforms on all understanding tasks with an average of 5.13% when compared to an important baseline, Yo’LLaVA(1.3B), demonstrating the potential of scaling UniCToken to achieve state-of-the-art performance.
We evaluate personalized generation capabilities of UniCTokens on Dreambench ruiz2023dreambooth and Yo’LLaVA Dataset nguyen2024yo . Our approach significantly enhances the capabilities of vanilla Show-o. Additionally, our performance surpasses that of Yo’Chameleon, particularly on CLIP-T, which measures the model’s ability for controllable generation. This improvement can be attributed to our more effective prompt design and mutual enhancement between personalized tasks. Notably, we also outperform the significant generative baseline, Dreambooth, with the same training data, demonstrating that our method can generate high-quality and realistic images containing user-provided concepts.
@@ 4.4 Ablation Study and Analysis
Stage Personalized Understanding Personalized Generation PARG Rec. VQA QA Pure Gen. People Gen. Weight BLEU Score BLEU Score CLIP - I CLIP-T DINO Face-Simi Score CLIP-I Stage 1 0.637 0.494 0.538 0.540 0.600 0.524 0.278 0.446 0.060 0.204 0.511 Stage 2 w/o 1 0.616 0.479 0.488 0.527 0.581 0.695 0.243 0.582 0.210 0.297 0.695 Stage 2 0.621 0.497 0.532 0.538 0.605 0.752 0.280 0.648 0.349 0.349 0.750 Stage 3 0.790 0.503 0.523 0.544 0.603 0.750 0.282 0.646 0.334 0.334 0.749
@@
Different Training Strategies.
Our approach utilizes a three-stage training strategy to facilitate information transfer across tasks. We then validate this design, beginning from generation, especially examining the inclusion of Stage 1. Omitting Stage 1 yields a variant of Text inversion distinguished by an increasing parameter count. Tab. 3 emphasizes the essential role of Stage 1, particularly in personalized attribute-reasoning generation, omitting Stage 1 leads to significant degradation in generation quality. This further corroborates the existence of information transfer across tasks in our design. Stage 3 does not update the generative parameters, resulting in a subtle influence on the generation task. In the context of the understanding task, we present the results of the evaluations conducted across all stages. Although performance in the understanding task declines in Stage 2 to facilitate generation, the subsequent injection of fine-grained visual information from Stage 3 enhances understanding capabilities. Results demonstrate that our approach achieves optimal performance, facilitating mutual enhancement between the two tasks.
@@ 5 Conclusion
We present UniCTokens, an innovative framework designed to personalize a unified VLM by training unified concept tokens. Our proposed three-stage training strategy enables UniCTokens to efficiently enhance personalized understanding and generation, facilitating cross-task information transfer to achieve mutual enhancement. Experimental results demonstrate that, while maintaining a smaller model size and fewer training samples, UniCTokens achieves state-of-the-art performance in concept understanding, concept generation, and attribute-reasoning generation tasks. The insights derived from our analysis are noteworthy, shedding light on the development of unified VLMs. Furthermore, the advancement of personalized AI holds significant promise for improving human lives by enabling tailored interactions, enhancing creativity, and offering solutions that align with individual needs and preferences, thereby fostering a more intuitive and responsive technological ecosystem.
6 Acknowledgments
This work is supported by the National Key R&D Program of China (2024YFA1014003), National Natural Science Foundation of China (92470121, 62402016), and High-performance Computing Platform of Peking University.
@@ Appendix A Additional Implement Detail
We summarize the full three-stage training setup in Tab. 4. Unless otherwise stated, before Stage 3 we run $k$-means with K=2K\!=\!2, using cosine distance ($d(\mathbf{u},\mathbf{v})=1-\tfrac{\mathbf{u}^{\top}\mathbf{v}}{\lVert\mathbf{u}\rVert\lVert\mathbf{v}\rVert}$). At each selection step, we retain the cluster containing the largest number of patches.
@@ Appendix B Additional Experiment Setup
For understanding, we evaluate the model on three tasks: personalized recognition, VQA, and QA. In the Recognition task, we compute the recall for both positive and negative samples and report their arithmetic mean. For VQA and QA, we evaluate the predicted answers using BLEU papineni2002bleu scores against ground truth, and further apply an LLM-based evaluation that scores responses from 0 to 1 based on alignment with key points in the reference answers. For general personalized image generation, we adopt prompts from the DreamBooth ruiz2023dreambooth dataset, and evaluate image quality using CLIP-based metrics: CLIP-I (image-to-image similarity) and CLIP-T (image-to-text similarity). Since half of our concepts are human subjects, we additionally employ the off-the-shelf ArcFace model deng2019arcface to measure facial similarity between generated and reference images. For the evaluated GPT-4o and GPT-4o used for scoring, we utilize different versions of them to make fair comparisons.
@@
  • Yo’LLaVA nguyen2024yo : One of the earliest work to explore VLM personalization. Following the original paper, we manually construct hard negative datasets and train Yo’LLaVA with different sizes of base model(Phi-1.5 abdin2024phi , 1.3B and Vicuna chiang2023vicuna , 13B) to make a fair comparison.
  • MC-LLaVA an2024mc : A model designed for enhancing multi-concept personalization tasks. Utilizing dual textual and visual prompts, it serves as a strong baseline for personalized understanding.
  • RAP-MLLM hao2024remember : We utilize the RAP-LLaVA model and follow the RAP-MLLM approach to construct a personalized database for each concept. To obtain the capability of understanding personalization, RAP-MLLM conducted a post-training on a dataset of 260K in size.
  • Dreambooth ruiz2023dreambooth : DreamBooth enhances the personalization capability of generative models by allowing users to fine-tune models with a limited number of images, thereby producing highly specific and context-aware outputs. For fair comparison, we utilize different number of training data (10, 3,000) to better evaluate Dreambooth.
  • Text Inversion gal2022image : Text inversion is a technique that transforms textual prompts into corresponding visual representations, enabling the generation of images based on description.
@@ Appendix C Catastrophe Forgetting
When a model acquires new knowledge, there exists the risk of catastrophic forgetting of previously learned information. Following Yo’Chameleon nguyen2025yochameleon , we evaluate the personalized model on several benchmarks assessing general capabilities. We conducted a comparative analysis against the original Show-o across well-established multimodal benchmarks: GenEval ghosh2023geneval , MMMU yue2024mmmu , and POPE li2023evaluating . The experimental results shown in Fig. 10 indicate that, despite training through a three-stage process and having task-specific tokens, model performance does not diminish after each stage, effectively preserving the model’s general capabilities.
@@
[…]
@@ Appendix D Additional Related Work
A multitude of efforts have been dedicated to employing a single model for both understanding and generation. SEED ge2023planting adapted image representation through discrete tokenization for language modeling, leveraging autoregressive conditioning for generation via an external decoder.  ge2024seed ; dong2023dreamllm restructured conditioning information aggregation but still relied on extra modules for generation. Chemeleon team2024chameleon is an early hybrid unified model capable of generating and understanding intert-leaved text-image content. Janus wu2024janus claims that understanding and generation require distinct information, employing different tokenizers for each task. Emu3 wang2024emu3 converts images, text, and video into discrete tokens, enabling joint training of a Transformer on multimodal sequences.  xie2024show ; zhou2024transfusion utilize a hybrid of autoregressive and diffusion methods for text and image processing. The above-mentioned work all focuses on general tasks, neglecting exploration in personalization scenarios. In this paper, we employ Show-o xie2024show as the backbone to efficiently achieve unified personalization without forgetting the general capability.
@@
The establishment of unified models seeks to optimize the strengths of understanding and generation, allowing information transfer between tasks and mutual improvement. Early efforts to link these tasks mainly utilized a serial processing paradigm.  luo2024llm ; liu2024synthvlm ; dunlap2023diversify employed image captioning models to generate textual conditions for text-to-image models. While these methods highlight the potential of leveraging understanding for generation, they lack end-to-end optimization and frequently suffer from information loss.  wu2024liquid ; tong2024metamorph ; fan2024fluid ; lin2025pixwizardversatileimagetoimagevisual ; fang2023instructseq explored deeper integration of understanding and generation. MetaQuery pan2025transfer utilizes learnable queries on frozen VLMs to extract conditions for generation, but it primarily emphasizes how understanding aids generation while neglecting the inverse. In this paper, we propose a three-stage training strategy that achieves information transfer and mutual enhancement between tasks in personalization scenarios, providing insights for the broader development of unified models.
Reasoning in Understanding and Generation.
The success of Deepseek-R1 guo2025deepseek promotes the exploration of reasoning. The recent work focuses on designing new reward zhang2024mavismathematicalvisualinstruction ; li2025chemvlm ; zhang2025critic , constructing valuable samples muennighoff2025s1 ; li2025imagine ; zhang2025scaling and building reasonable reinforcement learning algorithm zhang2025perlpermutationenhancedreinforcementlearning ; zhang2025pixelcraftmultiagenthighfidelityvisual ; wang2025vl to fully unleash the potential of LLMs wang2025emergent and VLMs xu2025visulogic ; li2025taco ; li2025miv . Many works wei2025tiif ; zhangmme tend to transfer the reasoning capability into image generation process.  guo2025can analysis the effectiveness of Chain-of-Thought(CoT) in generation, while  jiang2025t2i utilizing semantic- and token-level CoT to handle complex reasoning scenarios. Inspired by there work, we propose complex personalized attribute-reasoning generation, which better modeling the real world user query.
@@ Appendix E Additional Qualitative Results
Appendix F Additional Quantitative Experiments
@@
Learnable Token
Length
Und. Gen.
@@
Cost of Adding a New Concept.
Tab. 5 reports per-concept FLOPs and wall-clock time under a unified setup. Methods that personalize with large per-concept image sets—exemplified by Yo’Chameleon (about 1,100 images)—incur the highest cost ($7{\times}10^{17}$ FLOPs; 120 min), whereas generator-only tuning (DreamBooth) is far cheaper ($2{\times}10^{15}$ FLOPs; 7 min) but limited in scope. Among unified models, UniCTokens achieves a favorable cost–capability trade-off: $1.3{\times}10^{17}$ FLOPs and 25 min per concept, yielding lower time than Yo’Chameleon.
Type FLOPs / Concept Time / Concept
Yo’LLaVA-13B $1.5\times 10^{17}$ 30 min
Yo’LLaVA-Phi $1.6\times 10^{16}$ 10 min
DreamBooth (SD) $2\times 10^{15}$ 7 min
Yo’Chameleon $7\times 10^{17}$ 120 min
UniCTokens (ours) $1.3\times 10^{17}$ 25 min
Table 5: Per-concept training cost.
Training Scheme: Three-Stage vs. Joint.
Tab. 6 contrasts a single-stage joint baseline with our three-stage schedule. The staged schedule delivers consistent gains in personalized understanding and generation—with the largest improvements on personalized attribute-reasoning generation. These results indicate that staging is not redundant: an understanding warm-up first stabilizes concept bindings; generation bootstrapped from these bindings improves conditioning quality; the final stage feeds generation signals back to understanding. Consequently, cross-task transfer is strengthened without increasing model capacity.
Method Rec. VQA QA Pure Gen. People Gen. PARG
Weight GPT GPT CLIP-I CLIP-T Face-Sim. Score CLIP-I
Joint Training 0.709 0.489 0.565 0.681 0.258 0.288 0.269 0.678
3-Stage (Ours) 0.790 0.523 0.603 0.750 0.282 0.334 0.359 0.749
Table 6: Joint vs. three-stage. PARG = Personalized Attribute-Reasoning Generation.
Judge Models and Human Alignment.
Beyond GPT-based scoring and the classical metrics reported in the main paper, we additionally evaluate with Gemini-2.5-Pro and a human study on 300 samples (five per concept per task). As shown in Tab. 7, the relative ordering of methods is consistent between Gemini and human judges across VQA, QA, and personalized attribute-reasoning generation: UniCTokens ranks highest where reported. These trends support using LLM judges as a practical proxy while remaining aligned with human preferences.
Gemini-2.5-Pro
Methods VQA QA PARG
Yo’LLaVA–Phi 0.492 0.498
DreamBooth (SD) 0.066
Yo’Chameleon 0.489 0.495 0.279
UniCTokens (ours) 0.527 0.601 0.371
Human Study
Methods VQA QA PARG
Yo’LLaVA–Phi 0.679 0.622
DreamBooth (SD) 0.012
Yo’Chameleon 0.670 0.623 0.272
UniCTokens (ours) 0.700 0.683 0.352
Table 7: LLM and human scores on our bench. PARG = Personalized Attribute-Reasoning Generation.
@@ Appendix G Details of Dataset
The dataset’s data sources comprise animals and objects obtained from MC-LLaVA an2024mc , Yo’LLaVA nguyen2024yo and MyVLM alaluf2024myvlm , with images of individuals sourced from Yo’LLaVA and various online platforms. All images and generated training data have been subjected to rigorous human validation processes. We present a set of images along with extra textual information in Fig. 13.